# Federated Instruction Tuning of LLMs with Domain Coverage Augmentation

### v3

@@ A.2.1 Data Heterogeneity Is Not Matter In FedDIT
Following the traditional approach of constructing different degrees of non-iid, which are widely used in federated learning (Wang et al., 2020; Yurochkin et al., 2019), we adopt Dirichlet distribution to construct various heterogeneity and use k-means with the cluster num $\xi=100$ to pseudo labeling instructions. Dirichlet distribution is affected by the hyperparameter $\alpha$, which enhances heterogeneity with a smaller $\alpha$ and decreases heterogeneity with a larger $\alpha$. We choose four widely used heterogeneity, which are $\alpha=[0.01,0.1,1,10]$ (Ye et al., 2023; Zhang et al., 2023b; li2021model). Figure LABEL:fig:data_heterogeneity shows a visualization of the data distribution with different heterogeneity on the code dataset.
@@ A.6 Implementation Details
We consider FedDIT in the cross-device scenario, $N$ = 10 clients, $\mathcal{R}$ = 30 rounds, where we randomly sample 2 clients to be available for each round. Then, each available client performs FedDIT for 10 steps with AdamW optimizer, and the batch size is $B=32$ in a round. The initial learning rate is $55e-$ 5 with a cosine learning rate scheduler. Our experiment utilizes the widely used LLM, Llama3-8B [^5] https://huggingface.co/meta-llama/Meta-Llama-3-8B as the base model with 2048 max sequence length and adopts LoRA tuning method. The rank of LoRA is 16, and the scalar alpha is 16. For k-means (Vardakas & Likas, 2023), we set cluster num $\xi$ = 10 and for FedDCA we set the similarity threshold $\alpha$ = 0.7. For FedDCA, we set the temperature parameter $\tau$ = 0.5 for contrastive learning (Khosla et al., 2020). We utilize bge-large-en-v1.5666https://huggingface.co/BAAI/bge-large-en-v1.5 as both the client and server’s encoder as default, which outputs embeddings of 1024 dimensions. While we utilize all-MiniLM-L6-v2777https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 as the client’s small encoder which outputs embeddings of 384 dimensions and bge-large-en-v1.5 as the server’s encoder for FedDCA.

### v4

@@ A.2.1 Data Heterogeneity Is Not Matter In FedDIT
Following the traditional approach of constructing different degrees of non-iid, which are widely used in federated learning (Wang et al., 2020; Yurochkin et al., 2019), we adopt Dirichlet distribution to construct various heterogeneity and use k-means with the cluster num $\xi=100$ to pseudo labeling instructions. Dirichlet distribution is affected by the hyperparameter $\alpha$, which enhances heterogeneity with a smaller $\alpha$ and decreases heterogeneity with a larger $\alpha$. We choose four widely used heterogeneity, which are $\alpha=[0.01,0.1,1,10]$ (Ye et al., 2023; Zhang et al., 2023b; Li et al., 2021). Figure LABEL:fig:data_heterogeneity shows a visualization of the data distribution with different heterogeneity on the code dataset.
@@ A.6 Implementation Details
We consider FedDIT in the cross-device scenario, $N$ = 10 clients, $\mathcal{R}$ = 30 rounds, where we randomly sample 2 clients to be available for each round. Then, each available client performs FedDIT for 10 steps with AdamW optimizer, and the batch size is $B=32$ in a round. The initial learning rate is $55e-$ 5 with a cosine learning rate scheduler. Our experiment utilizes the widely used LLM, Llama3-8B [^5] https://huggingface.co/meta-llama/Meta-Llama-3-8B as the base model with 2048 max sequence length and adopts LoRA tuning method. The rank of LoRA is 16, and the scalar alpha is 16. For k-means (Wu, 2012), we set cluster num $\xi$ = 10 and for FedDCA we set the similarity threshold $\alpha$ = 0.7. For FedDCA, we set the temperature parameter $\tau$ = 0.5 for contrastive learning (Khosla et al., 2020). We utilize bge-large-en-v1.5666https://huggingface.co/BAAI/bge-large-en-v1.5 as both the client and server’s encoder as default, which outputs embeddings of 1024 dimensions. While we utilize all-MiniLM-L6-v2777https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 as the client’s small encoder which outputs embeddings of 384 dimensions and bge-large-en-v1.5 as the server’s encoder for FedDCA.